A Bibliography of Publications of Eric Grosse

Eric Grosse

Google

1600 Amphitheatre Parkway Mountain View, CA 94043

USA

Tel: none FAX: none

E-mail: eric@n2vi.com (Internet)

01 March 2018 Version 1.48

Abstract

This bibliography records publications of Eric Grosse.

Title word cross-reference

δ^{2} [BDG81].

-Formula [BDG81].

11th [USE02]. 15th [ARL78].

528 [FHS78, GG99].

77 [GG99].

9 [CGP^{+}02].

Access [BG95]. ACM [STO85]. adapting [GG99]. Additive [FGS83]. Aitken [BDG81]. Alabama [ARL78]. Algorithm [FHS78, Fox79, GG99]. Algorithms [CDG88, Gro90a]. Amorphous [PFG83b]. Analysis [ARL78, CCG^{+}76, CGR87, GW78, JG76]. Analyzer [CGR83]. Animation [CG90a, CG90b]. Annual [STO85]. Applications [YLS^{+}01]. Applied [BCG94, GN03]. Approximation [FGS83, Gro80, Gro81, Gro86, Gro90a, Gro93]. Arc [BG87]. Army [ARL78]. Arsenal [ARL78]. Aspects [CGR87]. Assessment [Boi97]. Asymptotic [BDG81]. Attenuated [BOG87]. August [USE02]. Authentication [GU13].

be [GM87]. Bifurcations [CGP92]. big [Gro91b]. Building [DGG^{+}08, GW78].

CA [USE02]. Calculation [PFG83b]. can [GM87]. Carlo [PFG83b]. Catalogue [Gro90a]. CAzM [CGR83]. Characteristics [CGP92]. Characterization [BOG87]. Charges [CFG89]. Circuit [CGR83, CGR87]. Circuit-Analyzer [CGR83]. Circuits [BCF^{+}85]. Circular [BG87]. Cloud [GHR^{+}10]. Coefficients [GH94]. Colors [Gro85]. Command [ARL78]. Comment [GG99]. Community [DGG^{+}08]. Computational [CDG88, CG91a, CGR87]. Computers [ARL78]. Computing [STO85, BBD95, CG89, CGP92, DGG^{+}08, GHR^{+}10]. Conference [ARL78, Boi97, GH92]. Conformal [BG87]. Connect [Gro91a]. Constants [GG99]. Control [BG95]. Coupled [BCG94]. Cryptographic [FGR92]. CSE [Gro96]. Current [CGP92]. Current-Voltage [CGP92].

Data [BBD95]. Decomposition [BCG94]. Density [Gro81, JG76]. Development [ARL78]. Device [CFG89]. Devices [BCF^{+}85, CGP92]. Digital [BBD95]. Diminishing [CGR86]. Display [CG91b]. Distributed [BDG^{+}95]. Distribution [DG85, DG87, PFG83a]. Domain [ADG87, BCG94]. Dynamic [CG91c].

Editor [CG83]. Electron [Gro81]. Electronic [DG87, DG85]. Electronically [DG89]. Enhancement [Boi97]. environments [GH92]. Equations [BCG94]. Expansions [BDG81]. Extracting [CFG89]. Extrapolation [BDG81].

Files [Gro92, Gro91b]. Filtered [BG95]. Finding [ADG87]. Fingered [BFG85]. Fitting [CDG88, CFG89]. Folds [CGP92]. Formula [BDG81]. Fortran [GG99]. Framework [FHS78, Fox79]. Francisco [USE02]. Functions [CG91b, PFG83a].

Gradient [CFG89]. Graphics [Gro78]. Grove [CG83]. Guide [CCG^{+}76].

Hand [BFG85]. Harmful [GM87]. Hearing [CG91c]. high [GH92]. high-level [GH92].

IFIP [Boi97, GH92]. Improved [GH94]. Independent [BDG^{+}95]. Inferno [Gro97]. Infrastructure [YLS^{+}01]. Interpolation [JG76]. Interpretation [JG76]. Ion [PFG83a]. Ions [PFG83b]. IPv4 [GN03]. IPv4/IPv6 [GN03]. IPv6 [GN03]. Island [STO85].

July [Boi97].

Karlsruhe [GH92]. Kingdon [Boi97].

Knots [GH94].

Language [GC83]. Languages

[CG79, Gro78]. Least [Gro89]. Level

[Gro85, GH92]. Library

[CCG^{+}76, CCGH80, FHS78, Fox79]. line [JG76]. Lists [FGR92]. Local [BG95, CDG88, CG91a, CGS91]. Location [BDG^{+}95]. Location-Independent [BDG^{+}95]. Loess [CG91c, Gro89].

Machine [GG99]. Macromodeling

[CGR83]. Mail [DG87, DG85]. Map [JG76]. Mapping [BG87]. Maps [Gro81]. Mathematical [ADG87, BDGR95, DG87, DG89, DG85]. May [STO85]. Membership [FGR92]. Methods [CDG88, CG91a]. Minimal [JG76]. Mirroring [Gro95]. Missile [ARL78]. Model [GW78]. Models [CGS91]. Modified [BDG81]. Monte [PFG83b]. Moving [Gro89]. Multi [BFG85]. Multi-Fingered [BFG85]. Multidimensional [FGS83]. Multilayer [BOG87]. Multivariate [Gro89].

NA-Net [DGG^{+}08]. Naming [BDG^{+}95]. NAPLUG [CCG^{+}76]. Net [DGG^{+}08]. Netlib [Gro90b, BDGR95, DGG^{+}08, Gro91b, Gro92]. Network [Gro96, GN03]. News [Gro90b, Gro92, Gro91b]. Non [BGR88]. Non-Obtuse [BGR88]. Nondestructive [BOG87]. Numerical [ARL78, Boi97, CCG^{+}76, CCGH80, CG79, GW78].

Obtuse [BGR88]. Optimization [Gro81]. Our [Gro91a]. Oxford [Boi97].

Parallel [BCG94]. Partitioned [YLS^{+}01]. Philosophy [CG89]. Pine [GC83]. Plan [CGP^{+}02]. Plots [Gro85]. Polygons [BGR88, BG87, CG91b]. Polynomials [JG76]. Portable [FHS78, Fox79]. Prehension [BFG85]. problem [GH92]. Proceedings [STO85, ARL78, Boi97, GH92, USE02]. Processors [GN03]. Program [CCG^{+}76]. Programming [GH92, Gro78, GC83, Gro96]. Properties [CDG88]. Protection [FGR92]. Protium [YLS^{+}01]. Providence [STO85]. Public [ADG87].

Quality [Boi97].

Ranges [PFG83a]. Real [Gro97]. Redstone [ARL78]. Refinement [GW78]. Reflection [BOG87]. Regression [CDG88, CG91a, CGS91]. Remark [Fox79]. Repositories [BBD95, BDG^{+}95]. Repository [Gro95, BDGR95]. Research [ARL78]. Resonant [BOG87]. Restyling [Gro78]. Rhode [STO85]. Ridge [JG76]. Ridge-line [JG76]. Rounding [GH94]. Roundtable [GHR^{+}10].

San [USE02]. Scale [GU13]. Scientific [BBD95, CG89, CG90a, CG90b, DGG^{+}08, GH92]. Search [Gro92]. Security [CGP^{+}02, USE02]. Seeing [CG91c]. Self [GG99]. Self-adapting [GG99]. Semiconductor [CGP92]. Seventeenth [STO85]. Shaded [CG91b]. Shall [Gro91a]. Shopping [DG89]. Silicon [BCF^{+}85].

Simulation [BCF^{+}85, CGR86, Gro93]. Simulations [CFG89]. Smoothing [Gro89]. Software [ADG87, BBD95, Boi97, BDG^{+}95, BDGR95, CG79, DG87, DG89, Gro78, Gro91a, DG85]. solving [GH92]. Sound [CG91b]. Space [CG91b]. Spanning [JG76]. Spectral [Gro86]. Spectroscopy [BOG87]. Spline [FGS83, Gro80, Gro86, GH94]. Splines [CGR86]. Squares [Gro89]. Stable [BFG85]. Structures [BOG87]. Support [BBD95, CCGH80]. Surfaces [CG91c]. Symposium [STO85, USE02].

Targets [PFG83b]. TC2 [Boi97, GH92]. TC2/WG [Boi97, GH92]. Techniques [CG90a, CG90b]. Tensor [Gro80]. Theory [STO85]. Three [CG91b]. Time [CG91b]. Tools [CG89, Gro91a]. Total [BOG87]. Transient [BCF^{+}85]. Transistor [CFG89]. Transition [GN03]. Transport [BCG94, PFG83b]. Trees [JG76]. Triangulation [BGR88].

U.S [ARL78]. Underflow [GM87]. United [Boi97]. User [CCG^{+}76]. Using [CG91b].

Variables [CG91b]. Variation [CGR86]. Vectorized [PFG83b]. Via [DG87, DG85]. video [CG90b]. Virtual [BDG^{+}95]. VLSI [Gro93]. Voltage [CGP92].

WG [Boi97, GH92]. Working [Boi97, GH92]. WWW [BG95].

Z [FHS78, Fox79].

References

Astfalk:1987:FPD

[ADG87] Greg Astfalk, Jack Dongarra, and Eric Grosse. Finding public domain mathematical software. Numerical Analysis Manuscript 87-5, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1987.

ARO:1978:PAN

[ARL78] Proceedings of the 1978 Army Numerical Analysis and Computers Conference (15th: 1978: U.S. Army Missile Research and Development Command, Redstone Arsenal, Alabama), number 78-3 in ARO report. U.S. Army Research Office, 1978.

Boisvert:1995:DSD

[BBD95] Ronald Boisvert, Shirley Browne, and Jack Dongarra. Digital software and data repositories for support of scientific computing. In Digital Libraries Forum. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, May

1995. URL ftp://netlib.bell-
labs.com/netlib/srwn/srwn09.

ps.gz. McLean, Virginia.

Bank:1985:TSS

[BCF^{+}85] R. E. Bank, W. M. Coughran, Jr., W. Fichtner, E. H. Grosse, D. J. Rose, and R. K. Smith. Transient simulation of silicon devices and circuits. IEEE Trans. on Computer-Aided Design, CAD-4: 436–451, 1985. (also IEEE Trans. on Electron Devices ED-32).

Bjorstad:1994:PDD

[BCG94] Petter Bjørstad, W. M. Coughran, Jr., and Eric Grosse. Parallel domain decomposition applied to coupled transport equations. In David E. Keys and Jinchao Xu, editors, Domain Decomposition Methods in Scientific and Engineering Computing, pages 369–380. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-5171-3. LCCN QA402.2 I55 1993.

URL ftp://cm.bell-labs.com/
cm/cs/doc/94/4-03.ps.gz.

Bjorstad:1981:EAE

[BDG81] Petter E. Bjørstad, Germund Dahlquist, and Eric H. Grosse. Extrapolation of asymptotic expansions by a modified Aitken δ^{2}- formula. BIT, 21(1):56–65, 1981. CODEN BITTEL, NBITAB. ISSN 0006-3835 (print), 1572-9125 (electronic).

Browne:1995:LIN

[BDG^{+}95] Shirley Browne, Jack Dongarra, Stan Green, Keith Moore, Theresa Pepin, Tom Rowan, Reed Wade, and Eric Grosse. Locationindependent naming for virtual distributed software repositories. In Symposium on Software Reusability. ACM Press, New York, NY 10036, USA, April

1995. URL ftp://netlib.bell-
labs.com/netlib/srwn/srwn07.

ps.gz. Seattle, Washington.

Browne:1995:NMS

[BDGR95] Shirley Browne, Jack Dongarra, Eric Grosse, and Tom Rowan. The Netlib Mathematical Software Repository. D-Lib magazine: the magazine of the Digital Library Forum, September 1995. ISSN 1082-9873. URL http://www.

cnri.reston.va.us/home/dlib.
html.

Baker:1985:SPM

[BFG85] Brenda S. Baker, S. J. Fortune, and Eric H. Grosse. Stable prehension with a multi-fingered hand. In STOC’85 [STO85], pages 114–120. ISBN 0-89791-151-2. LCCN QA 76.6 A13 1985.

Bjorstad:1987:CMC

[BG87] P. E. Bjørstad and E. H. Grosse.

Conformal mapping of circular arc polygons. SIAM J. Sci. Stat. Comput., 8:19–32, 1987. CODEN SI-JCD4. ISSN 0196-5204.

Baker:1995:LCF

[BG95] Brenda S. Baker and Eric Grosse.

Local control over filtered WWW access. In Fourth International World Wide Web Conference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 1995.

URL http://www.w3.org/pub/
Conferences/WWW4/Papers/117/

. Boston, MA.

Baker:1988:NOT

[BGR88] Brenda Baker, Eric Grosse, and Conor Rafferty. Non-obtuse triangulation of polygons. J. Discrete and Computational Geometry, 3:147–168, 1988. URL http:/

/cm.bell-labs.com/cm/cs/doc/
85/nonobtuse.pdf.

Bosacchi:1987:NCM

[BOG87] Bruno Bosacchi, Robert C. Oehrle, and Eric Grosse. Nondestructive characterization of multilayer structures by resonant attenuated total reflection spectroscopy. Applied Physics Letters, 51:158–160, 1987.

Boisvert:1997:QNS

[Boi97] Ronald F. Boisvert, editor. The Quality of Numerical Software: Assessment and Enhancement: Proceedings of the IFIP TC2/WG 2.5 Working Conference on the Quality of Numerical Software, Oxford, United Kingdon, 8–12 July 1996. Chapman Hall on behalf of IFIP, London, 1997. ISBN 0-412-80530-8.

Chan:1976:NAP

[CCG^{+}76] T. F. Chan, W. M. Coughran, Jr., E. H. Grosse, M. T. Heath, and F. T. Luk. Numerical analysis program library user’s guide (NAPLUG). User Note 82, SLAC Computing Services, 1976.

Chan:1980:NLS

[CCGH80] Tony F. Chan, William M. Coughran, Jr., Eric H. Grosse, and Michael T. Heath. A numerical library and its support. ACM Trans. Math. Software, 6(2):135– 145, June 1980. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Cleveland:1988:RLF

[CDG88] William S. Cleveland, Susan J. Devlin, and Eric Grosse. Regression by local fitting: Methods, properties, and computational algorithms. J. Econometrics, 37:87– 114, 1988.

Coughran:1989:ETC

[CFG89] W. M. Coughran, Jr., W. Fichtner, and Eric Grosse. Extracting transistor charges from device simulations by gradient fitting. IEEE Trans. on Computer-Aided Design, 8:380–394, 1989.

Coughran:1979:NLN

[CG79] W. M. Coughran, Jr. and E. H.

Grosse. New languages for numerical software. SIGNUM Newsletter, 14:73–75, 1979. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Coughran:1983:GE

[CG83] W. M. Coughran, Jr. and E. H.

Grosse. The grove editor. Numerical Analysis Manuscript 83-3, ATT Bell Laboratories, 1983.

Coughran:1989:PSC

[CG89] W. M. Coughran, Jr. and Eric Grosse. A philosophy for scientific computing tools. SIGNUM Newsletter, 24(2/3):2–9, April/ July 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Coughran:1990:TSAa

[CG90a] W. M. Coughran, Jr. and Eric Grosse. Techniques for scientific animation. SPIE Proceedings, 1259:72–79, 1990.

Coughran:1990:TSAb

[CG90b] W. M. Coughran, Jr. and Eric Grosse. Techniques for scientific animation—video. SPIE Proceedings, 1259-V:22:00–35:28, 1990. ISBN 0-8194-0306-7.

Cleveland:1991:CML

[CG91a] William S. Cleveland and Eric Grosse. Computational methods for local regression. Statistics and Computing, 1(1):47–62, 1991. CO-DEN STACE3. ISSN 0960-3174 (print), 1573-1375 (electronic).

URL ftp://cm.bell-labs.com/
cm/cs/doc/91/4-04.ps.gz.

Coughran:1991:DFT

[CG91b] W. M. Coughran, Jr. and Eric Grosse. Display of functions of three space variables and time using shaded polygons and sound. In Gaffney and Houstis [GH92], pages 271–276. ISBN 0-444-89176-5. LCCN QA76.6 .I1782 1992.

URL ftp://cm.bell-labs.com/
cm/cs/doc/91/4-09.ps.gz.

Coughran:1991:SHD

[CG91c] W. M. Coughran, Jr. and Eric Grosse. Seeing and hearing dynamic loess surfaces. In Interface’91 Proceedings, pages 224– 228. Springer-Verlag, Berlin, Heidelberg, New York, Tokyo, 1991.

URL ftp://cm.bell-labs.com/
cm/cs/doc/91/4-07.ps.gz.

Coughran:1992:CFB

[CGP92] W. M. Coughran, Jr., E. H. Grosse, and M. R. Pinto. Computing folds and bifurcations in current-voltage characteristics of semiconductor devices. In Workshop on Numerical Modeling of Processes and Devices for Integrated Circuits: NUPAD IV. Technical Digest, pages 149–153. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992.

Cox:2002:SP9

[CGP^{+}02] Russ Cox, Eric Grosse, Rob Pike, Dave Presotto, and Sean Quinlan. Security in Plan 9. In USENIX [USE02], pages 3–16. ISBN 1-931971-00-5. LCCN ???? URL http://

www.usenix.org/publications/
library/proceedings/sec02/cox.
html.

Coughran:1983:CCA

[CGR83] W. M. Coughran, Jr., E. H.

Grosse, and D. J. Rose. CAzM: A circuit-analyzer with macromodeling. IEEE Trans. on Electron Devices, ED-30:1207–1213, 1983.

Coughran:1986:VDS

[CGR86] William M. Coughran, Jr., Eric Grosse, and Donald J. Rose. Variation diminishing splines in simulation. SIAM J. Sci. Stat. Comput., 7:696–705, 1986. CODEN SI-JCD4. ISSN 0196-5204.

Coughran:1987:ACC

[CGR87] W. M. Coughran, Jr., Eric Grosse, and Donald J. Rose. Aspects of computational circuit analysis. In W. Fichtner and M. Morf, editors, VLSI CAD Tools and Applications, pages 105–127 (of x + 552). Kluwer Academic Press, Dordrecht, The Netherlands, 1987.

ISBN 0-89838-193-2. LCCN TK7874 .V5572 1987.

Cleveland:1991:LRM

[CGS91] William S. Cleveland, Eric Grosse, and William M. Shyu. Local regression models. In John M. Chambers and Trevor J. Hastie, editors, Statistical Models in S, pages 309–376 (of xv + 608). Wadsworth and Brooks/Cole, Belmont, CA, USA and Pacific Grove, CA, USA, 1991. ISBN 0-534-16765-9. LCCN QA276.4 .S65 1991.

Dongarra:1985:DMS

[DG85] J. J. Dongarra and E. Grosse. Distribution of mathematical software via electronic mail. SIGNUM Newsletter, 20(3):45–47, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1987:DMS

[DG87] Jack J. Dongarra and Eric Grosse. Distribution of mathematical software via electronic mail. Communications of the ACM, 30:403–407, 1987.

Dongarra:1989:SMS

[DG89] Jack Dongarra and Eric Grosse. Shopping for mathematical software electronically. IEEE Potentials, 8:37–38, February 1989. CODEN IEPTDF. ISSN 0278-6648 (print), 1558-1772 (electronic). condensed version of CACM paper.

Dongarra:2008:NNB

[DGG^{+}08] Jack Dongarra, Gene H. Golub, Eric Grosse, Cleve Moler, and Keith Moore. Netlib and NA-Net: Building a scientific computing community. IEEE Annals of the History of Computing, 30(2):30–41, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http:/

/ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=4544554.

Feigenbaum:1992:CPM

[FGR92] Joan Feigenbaum, Eric Grosse, and James A. Reeds. Cryptographic protection of membership lists. Newsletter of the International Association for Cryptologic Research, 9(1):16–20, 1992.

URL ftp://cm.bell-labs.com/
cm/cs/doc/91/4-12.ps.gz.

Friedman:1983:MAS

[FGS83] J. H. Friedman, E. H. Grosse, and W. Stuetzle. Multidimensional additive spline approximation. SIAM J. Sci. Stat. Comput., 4:291–301, 1983. CODEN SIJCD4. ISSN 0196-5204.

Fox:1978:AFP

[FHS78] P. A. Fox, A. D. Hall, and N. L.

Schryer. Algorithm 528: Framework for a portable library [Z]. ACM Trans. Math. Software, 4 (2):177–188, June 1978. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remarks [Fox79, GG99].

Fox:1979:RFP

[Fox79] Phyllis Fox. Remark on “Algorithm 528: Framework for a portable library [Z]”. ACM Trans. Math. Software, 5(4):524, December 1979. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [FHS78].

Grosse:1983:PPL

[GC83] E. H. Grosse and W. M. Coughran, Jr. The pine programming language. Numerical Analysis Manuscript 83-4, ATT Bell Laboratories, 1983. URL ftp:/

/cm.bell-labs.com/cm/cs/doc/
92/pine.ps.gz.

Gay:1999:SAF

[GG99] David M. Gay and Eric Grosse. Self-adapting Fortran 77 machine constants: Comment on Algorithm 528. ACM Trans. Math. Software, 25(1):123–126, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cm.

bell-labs.com/who/ehg/mach/
d1mach.ps;   http://www.acm.
org/pubs/citations/journals/
toms/cgi-bin/TOMSbibget?Gay:
1999:SAF;   http://www.acm.
org/pubs/citations/journals/
toms/cgi-bin/TOMScitation?Fox:
1978:AFP;   http://www.acm.
org:80/pubs/citations/journals/
toms/1999-25-1/p123-gay/. See

[FHS78].

Gaffney:1992:PEH

[GH92] P. W. Gaffney and E. N. Houstis, editors. Programming environments for high-level scientific problem solving: Proceedings of the IFIP TC2/WG 2.5 Working Conference on Programming Environments for High-Level Scientific Problem Solving, Karlsruhe. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-89176-5. LCCN QA76.6 .I1782 1992.

Grosse:1994:IRS

[GH94] Eric Grosse and John D. Hobby.

Improved rounding for spline coefficients and knots. Math. Comp., 63(207):175–194, 1994. CO-DEN MCMPAF. ISSN 0025-5718 (paper), 1088-6842 (electronic).

URL ftp://cm.bell-labs.com/
cm/cs/doc/93/4-13.ps.gz.

Grosse:2010:CCR

[GHR^{+}10] Eric Grosse, John Howie, James Ransome, Jim Reavis, and Steve Schmidt. Cloud computing roundtable. IEEE Security & Privacy, 8(6):17–23, November/ December 2010. CODEN ???? ISSN 1540-7993 (print), 1558-4046 (electronic).

Grosse:1987:UCB

[GM87] Eric Grosse and Cleve Moler. Underflow can be harmful. SIAM News, 20(6):1, 1987. ISSN 0036-1437.

Grosse:2003:NPA

[GN03] Eric Grosse and Lakshman Y.

N. Network processors applied to IPv4/IPv6 transition. IEEE Network, 17(4):35–39, 2003. CODEN IENEET. ISSN 0890-8044 (print), 1558-156x (electronic).

Grosse:1978:SRG

[Gro78] E. H. Grosse. Software restyling in graphics and programming languages. In ARL-78-3 [ARL78], pages 79–108.

Grosse:1980:TSA

[Gro80] E. H. Grosse. Tensor spline approximation. Linear Algebra and Its Applications, 34:29–41, 1980.

Grosse:1981:AOE

[Gro81] E. H. Grosse. Approximation and Optimization of Electron Density Maps. PhD thesis, Stanford University Computer Science Department, 1981. STAN-CS-80-835.

Grosse:1985:CLP

[Gro85] Eric Grosse. Colors for level plots. Numerical Analysis Manuscript 85-1, ATT Bell Laboratories, 1985.

                                        URL ftp://cm.bell-labs.com/
                                        cm/cs/doc/85/4-01.ps.gz.

Grosse:1986:SSA

[Gro86] Eric Grosse. Spectral spline approximation. In C. K. Chui, L. L. Schumaker, and J. D. Ward, editors, Approximation Theory V, pages 363–366 (of xviii + 654). Academic Press, New York, NY, USA, 1986. ISBN 0-12-174581-3. LCCN QA221 .I56 1986.

Grosse:1989:LMS

[Gro89] Eric Grosse. LOESS: Multivariate smoothing by moving least squares. In C. K. Chui, L. L. Schumaker, and J. D. Ward, editors, Approximation Theory VI, pages 299–302. Academic Press, New York, NY, USA, 1989. ISBN 0-12-174587-2. LCCN QA221 .I56 1989.

Grosse:1990:CAA

[Gro90a] Eric Grosse. A catalogue of algorithms for approximation. In J. Mason and M. Cox, editors, Algorithms for Approximation II, pages 479–514 (of 514). Chapman and Hall, London, England, 1990. ISBN 0-412-34580-3. LCCN QA221 .I54 1988.

URL ftp://cm.bell-labs.com/
cm/cs/doc/catalog.ps.gz.

Grosse:1990:NN

[Gro90b] Eric Grosse. Netlib news.

SIGNUM Newsletter, 1990. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

URL ftp://netlib.bell-labs.
com/netlib/news/index.html.

gz. Oct 90, Jan 91, Apr 92.

Grosse:1991:HSW

[Gro91a] Eric Grosse. How shall we connect our software tools. In Visualization’91 Proceedings, pages 292–296 (of xi + 437). IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2245-8. LCCN Q174 .V56

1991. URL ftp://cm.bell-labs.
com/cm/cs/doc/91/4-08.ps.gz.

Grosse:1991:NNB

[Gro91b] Eric Grosse. Netlib news: big files.

SIGNUM Newsletter, 26(4):4–5, October 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Grosse:1992:NNS

[Gro92] Eric Grosse. Netlib news: Search for files. SIGNUM Newsletter, 27(3):2–4, July 1992. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Grosse:1993:AVS

[Gro93] Eric Grosse. Approximation in VLSI simulation. Numerical Algorithms, 5:591–601, 1993. CO-DEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

                                        URL ftp://cm.bell-labs.com/
                                        cm/cs/doc/93/4-05.ps.gz.

Grosse:1995:RM

[Gro95] Eric Grosse. Repository mirroring. ACM Trans. Math. Software, 21(1):89–97, March 1995. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

                                        URL ftp://netlib.bell-labs.
                                        com/netlib/crc/mirror.ps.gz.

Grosse:1996:NPC

[Gro96] Eric Grosse. Network programming and CSE. IEEE Computational Science & Engineering, 3 (2):40–41, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). Discusses the emergence of systems like Java and Inferno for scientific visualization.

Grosse:1997:RI

[Gro97] Eric Grosse. Real Inferno. In Boisvert [Boi97], pages 270–279. ISBN 0-412-80530-8. URL ftp:

                                        //cm.bell-labs.com/inferno/
                                        real.ps.

Grosse:2013:AS

[GU13] Eric Grosse and Mayank Upadhyay. Authentication at scale. IEEE Security & Privacy, 11(1): 15–22, January/February 2013. ISSN 1540-7993 (print), 1558-4046 (electronic). URL http://www.

computer.org/cms/Computer.org/
ComputingNow/pdfs/AuthenticationAtScale.
pdf.

Grosse:1978:NAM

[GW78] E. H. Grosse and M. H. Wright.

Numerical analysis for model building and refinement. In National Resource for Computation Chemistry workshop, 1978.

Johnson:1976:IPM

[JG76] C. Johnson and E. H. Grosse. Interpolation polynomials, minimal spanning trees and ridge-line analysis in density map interpretation. American Crystallographic Association Program and Abstracts, 4 (2):48, 1976. CODEN ACRABY. ISSN 0569-4221.

Petersen:1983:DFI

[PFG83a] Wesley Petersen, Wolfgang Fichtner, and Eric Grosse. Distribution functions for ion ranges. Numerical analysis manuscript, ATT Bell Laboratories, 1983.

Petersen:1983:VMC

[PFG83b] Wesley Petersen, Wolfgang Fichtner, and Eric Grosse. Vectorized Monte Carlo calculation for the transport of ions in amorphous targets. IEEE Trans. on Electron Devices, ED-30:1011–1017, 1983.

ACM:1985:PSA

[STO85] Proceedings of the Seventeenth Annual ACM Symposium on Theory of Computing, Providence, Rhode Island, May 6–8, 1985. ACM Press, New York, NY 10036, USA, 1985. ISBN 0-89791-151-2. LCCN QA 76.6 A13 1985.

USENIX:2002:PUS

[USE02] USENIX, editor. Proceedings of the 11th USENIX Security Symposium 2002, August 5–9, 2002, San Francisco CA. USENIX, Berkeley, CA, USA, 2002. ISBN 1-931971-00-5. LCCN ????

Young:2001:PIP

[YLS^{+}01] Cliff Young, Y. N. Lakshman, Tom Szymanski, John Reppy, David Presotto, Rob Pike, Girija Narlikar, Sape Mullender, and Eric Grosse. Protium, and infrastructure for partitioned applications. In Eighth IEEE Workshop on Hot Topics in Operating Systems (HotOS-VIII). May 20–23, 2001, Schoss Elmau, Germany, pages 41–46. IEEE, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1040-X. US$135.00. URL

http://computer.org/CSPRESS/
CATALOG/pr01040.htm;   http:
//i30www.ira.uka.de/conferences/

HotOS/. IEEE catalog number PR01040.